What is the molal concentration of 30% ethanol solution (C2H5O5)?
1
Expert's answer
2020-10-06T09:19:31-0400

Assuming that 30% is a volume concentration (V/V), there are 300 mL of ethanol per 700 mL of water. Also, assuming the density of water equal to 1 g/mL, the volume of ethanol (solute) per 1 kg of water (solvent) can be determined:


"V(C_2H_5OH)=\frac{300mL\times1000g}{700g}=428.6mL"


The density of ethanol equals 0.789 g/mL.

Now, moles of ethanol per 1 kg of water can be calculated:


"n(C_2H_5OH)=\frac{m}{M}=\frac{\rho\times{V}}{M}=\frac{0.789g/mL\times428.6mL}{46.07g/mol}=7.34mol"


Therefore, the molal concentration is 7.34 m.

Answer: 7.34 m